OpenStudy (cookie_2046):

Solve 3(x - 2) + 5x = 18 A. all real numbers B. 1 C. 2 D. 3

OpenStudy (dehelloo):

https://www.youtube.com/watch?v=mK7Jj6HtevE yeeeeeeeeeeeeeeeeeew

OpenStudy (dehelloo):

straaaaaaaaaaya

OpenStudy (cookie_2046):

and how is that suppose to help me?

OpenStudy (anonymous):

x=3

OpenStudy (cookie_2046):

thxs

OpenStudy (anonymous):

Welcome :)

OpenStudy (adilalvi):

Solve 3(x - 2) + 5x = 18 3x-6+5x=18 18x-6=18 18x=18+6 18x=24 x=24/18 x=1.33